How many liters of a 60% solution should be added to 50 liters of a 30% solution to obtain a 40% solution?
Answer by Maths68(1474)

Soluton A
Amount = x
Concentration = 60%=0.6
================================================================================
Solution B
Amount = 50 liters
Concentration = 30%=0.3
================================================================================
Resultant Solution
Amount=50+x
Concentration=40%=0.4
================================================================================
(Amount of A * Concentration of A)+(Amount of B * Concentration of B) = (Amount of Resultant * Concentration of Resultant)
(x*0.6)+(50*0.3)=(50+x)*0.40
0.6x+15=20+0.4x
0.6x-0.4x=20-15
0.2x=5
0.2x/0.2=5/0.2
x=50
================================================================================
50 liters of a 60% solution should be added to 50 liters of a 30% solution to obtain a 40% solution?